# Euclidean Distance Matrix Completion via Asymmetric Projected Gradient Descent

### v1

@@ II-A Main Result
By our setting of parameters, (15) is a direct corollary of Lemma II.1 and Lemma II.2. The interval on $\eta^{\prime}$ comes from calculating $\frac{\eta\sigma_{r}^{\star}}{2}$ with the step size $\eta$ determined in Appendix B-A. What remains is to count iteration complexity using (15). Suppose choosing the largest possible $\eta^{\prime}=\frac{1}{336c_{I}\mu r}$, we need to show $(1-\frac{1}{336c_{I}\mu r})^{k_{0}-1}\leq\varsigma$, it requires
$k_{0}\geq\log(1/\varsigma)/\log(\frac{336c_{I}\mu r}{336c_{I}\mu r-1})+1,$
@@
Remark II.1.
It is worth noting that the requirement on $p$ in Lemma II.2, i.e., the establishment of regularity condition for the pseudo gradient, is of the same order in $\kappa$, $\mu$, $r$, and $n$ as its counterpart in classic LRMC theory[35, Lemma 3]. This observation leads us to conjecture that our techniques might be applied to tighten the estimate by Tasissa and Lai[20]. Meanwhile, the sample complexity for the OS-MDS is slightly suboptimal in $r$ compared with the best-known bound[65, Thm. 3.23]. The direct rank $d+2$ truncation of an EDM (i.e., the SVD-MDS) has been shown to be minimax sub-optimal in [49, Sec. IV]. Since the OS-MDS shares almost same empirical phase transition edge as the SVD-MDS[1], it suggests that the idea of using $\|\frac{1}{p}\mathcal{P}_{\Omega}\mathbf{D}^{\star}-\mathbf{D}^{\star}\|$ to construct spectral estimator might be overly conservative in practice.
III Numerical Experiments
The algorithm tested here is slightly different from the theoretical routine listed in Theorem II.1 (cf. Algorithm 2), in which BB stepsize without line search is employed, as inherited from[62, Ch. 2] [^9] The lower safeguard on BB stepsize is removed in most experiments, i.e., $\eta_{k}$ is allowed to be negative. Compared with pure GD, we found such sporadically occurring gradient ascent can numerically improve overall global convergence in all tested scenarios (especially the protein tests in Section III-E). The upper safeguard is retained to ensure algorithmic stability when $c_{IP}$ is large, c.f., (17).. The performance of incoherent projection-free OS-MDS, OS-MDS initialized GD refinement on s-stress (denoted as OS-MDS-GD) is also reported in Fig. 1 for comparison, more details on them can be found in[1].
Algorithm 2 APGD With BB Stepsize
1:  $\hat{\mathbf{P}}\hat{\mathbf{P}}^{T}=\mathcal{T}_{r}\left[-\frac{1}{2p}\mathbf{J}(\mathcal{P}_{\Omega}\mathbf{D}^{\star})\mathbf{J}\right],\,\mathbf{P}^{0}=\mathcal{P}_{I}(\hat{\mathbf{P}})$
@@ III Numerical Experiments
III-A Implementation Details
where Lemma C.6 guarantees the approximation behavior of $\hat{\sigma}_{1}^{\star}$, and $c_{1}$, $c_{2}$ are the parameters to tune. Since the APGD is more analogous to an analytical process rather than an algorithm used for benchmarking, we assume throughout all experiments that $\sigma_{1}^{\star}$ and $\mu$ are perfectly known. Unless otherwise specified, the ground truth point set is generated by a standard Gaussian distribution with embedding dimension $r=2$, and results presented here are obtained under $\eta_{\max}=10$. Algorithm 2 is terminated once $\|\mathbf{p}g^{k}\|_{F}\leq 1\times 10^{-6}$ or the iteration number reaches $N=1000$. In contrast, both the upper and lower safeguards, $\eta_{\max}=\frac{2}{p\sigma_{1}^{\star}\mu r\kappa}$ and $\eta_{\min}=\frac{1}{2p\sigma_{1}^{\star}\mu r\kappa}$, are adopted in Section III-D for better illustration of linear convergence. Please note that the calculation of $\mathbf{JAJ}$ for $\mathbf{A}\in\mathbb{R}^{n\times n}$ can be conducted in $\mathcal{O}(n^{2})$ flops, as it amounts to sequentially removing the mean value of $\mathbf{A}$’s columns and rows.
@@ III-A Implementation Details
where $\mathrm{g}_{2}^{k}$ can be viewed as the norm of the second term in the gradient of (4a), it cannot be computed unless direct access to $\mathbf{G}^{\star}$ is available. We conjecture that the limited numerical performance of APGD in the sample-limited regime can be attributed to at least three factors [^10] The second factor is believed to be the most dominant.: (i) the weakening of implicit regularization; (ii) the large residual between the pseudo-gradient and its population-level counterpart; (iii) the inconsistency between pseudo-gradient and the true gradient of (4a). These factors will be illustrated sequentially in the following sections.
Figure 1: The phase transition of regularization free OS-MDS and OS-MDS-GD when varying $p$ and $n$[1]. We claim a success if the spectral error falls below $1$ in (a), or the EDM recover rate is smaller than $10^{-3}$ in (b). For each $(n,p)$, the result is obtained by 100 independent trials.
Figure 2: (a) The phase transition of APGD when varying Bernoulli sample rate $p$ and the number of points $n$. The result is obtained by 100 independent trials. To verify Theorem II.1, we claim a success if the EDM recover rate falls below $10^{-3}$. In (b), (c), and (d), we vary the constant $c_{IP}$ in (17) when performing incoherent projection, while all other parameters remain unchanged. Subplot (e), (f), and (g) show selected phase transition curves when $n=500,1000,1500$ with $c_{IP}$ varies.
@@
TABLE I: Parameters of Test Proteins
@@
Figure 4: Trajectories of all three quotas in (16) while fixing $n=1500$. The result is obtained from 100 independent trials. (a), (d) shows the tendency of $r^{k}$ in selected succeed instances (96% when $p=0.1$, 47% when $p=0.05$). (b), (c) plots the corresponding $g_{1}^{k}$, $g_{2}^{k}$ when $p=0.1$. (e), (f) contains the records of $g_{1}^{k}$ and $g_{2}^{k}$ of selected failed instances (53% when $p=0.05$). These trajectories can be divided into two groups (17% in region 1, 36% in region 2).
III-B Phase Transition Under Gaussian Point Set
in (14) with some $c_{IP}>1$, the performance of APGD on Gaussian random point sets will be degenerated, as suggested by Fig. 2(b), (c), and (d). This observation indicates that the power of implicit regularization to maintain incoherence is weakened in the APGD case. More delicate evaluations are carried out in Fig. 2(e), (f), and (g). This reveals that APGD can still converge when the trimming step is removed, yet larger $c_{IP}$ gives rise to more oscillations and non-monotonicity in the transition curve, ultimately leading to broader transition interval measured by $p$.
By comparing Fig. 1(a) and Fig. 2(a), we find that the phase transition edge of APGD matches that of OS-MDS. This phenomenon confirms the prediction in Lemma II.2, and yet is quite unusual, as both non-convex Gaussian ensemble phase retrieval and structure-less LRMC do not necessarily rely on the spectral initialization to succeed[66][67]. Since the proof roadmap in[67] indicates that the success of rescaled eigenspace alignment initialization relies on the implicit incoherent regularization ability of the original fourth-order polynomial-type cost function for LRMC, we infer that when this property is compromised, the algorithm may start to behave selective on the quality of starting point. Alternatively, more dedicated initialization for APGD can somewhat improve the transition edge. However, the OS-MDS-GD strategy exhibits empirical convergence when started inside a slack region around the ground truth as depicted in Fig. 1(b), even without any incoherence projection.
@@ III-C Phase Transition Under Random Perturbation
where $\mathbf{E}_{n}$ is a Gaussian random matrix with $[\mathbf{E}_{n}]_{ij}\sim\mathcal{N}(0,1)$ and $\mathbf{P}^{\star}$ is generated from a uniform distribution over the hypercube $[-0.5,0.5]^{r}$. The noise level $\sigma_{n}$ is varied from $5\times 10^{-5}$ to $10$. The iteration will be terminated if $\|\mathbf{p}_{g}^{k}\|_{F}\leq 10^{-8}$, while all other parameters follow the same setup as in Fig. 2(b). Since the perturbation can be relatively small, we claim a success if $\mathrm{RE}\leq 10^{-5}$, and the resulting phase transitions are plotted in Fig. 4 for $n=500,1000,1500$, $r=2$, where $\|\Delta_{0}\|_{F}^{2}$ refers to $\mathrm{dist}(\mathbf{P}_{0},\mathbf{P}^{\star})^{2}$. Around $\sigma_{n}\leq 0.1$, i.e., inside the region where $\|\Delta_{0}\|_{F}^{2}\leq\sigma_{r}^{\star}$ according to Fig. 2(d), these plots depict that the convergence of APGD becomes sensitive primarily to the sample complexity. As $p\to 1$, the ^$\hat{\nabla}f(\mathbf{P})$ reduces to the gradient direction of the vanilla matrix factorization problem, which is known to exhibit local restricted strong convexity[52, Thm. 5]. Therefore, Fig. 3 also demonstrates the lower bound on $p$ necessary for this approximation to hold.
However, this dependence on $p$ is sub-optimal compared to the performance achieved by either OS-MDS-GD or trace minimization methods[20, Sec. IV][31, Sec. 5], indicating that the stabilization of pseudo gradient [^11] More specifically, ensuring the upper bound on Eq. (23) in Appendix B-B. needs more random samples to guarantee, even though Lemma II.2 predicts that the establishment of regularity condition should require the same order of sample complexity as in classical LRMC problems. Moreover, it is noteworthy that OS-MDS cannot achieve $\|\Delta_{0}\|_{F}^{2}\leq\sigma_{r}^{\star}$ with high probability at the sample configurations $\{(n,p):(500,0.08),(1000,0.05),(1500,0.04)\}$, which is why we allow Algorithm 2 to take negative stepsize, thereby facilitating global optimization.

### v2

@@ II-A Main Result
By our setting of parameters, (17) is a direct corollary of Lemma II.1 and Lemma II.2. The interval on $\eta^{\prime}$ comes from calculating $\frac{\eta\sigma_{r}^{\star}}{2}$ with the step size $\eta$ determined in Appendix B-A. What remains is to count iteration complexity using (17). Suppose choosing the largest possible $\eta^{\prime}=\frac{1}{336c_{I}\mu r\kappa^{2}}$, we need to show $(1-\frac{1}{336c_{I}\mu r\kappa^{2}})^{k_{0}-1}\leq\varsigma$, it requires $k_{0}\geq\mathcal{O}(336c_{I}\mu r\kappa^{2}\log(1/\varsigma))$, concluding the proof.∎
Remark II.1.
It is worth noting that the requirement on $p$ in Lemma II.2, i.e., the establishment of regularity condition for the pseudo gradient, is of the same order in $\kappa$, $\mu$, $r$, and $n$ as its counterpart in classic LRMC theory[38, Lemma 3]. This observation leads us to conjecture that our techniques might be applied to tighten the estimate by Tasissa and Lai[20]. Meanwhile, the sample complexity for the OS-MDS is slightly suboptimal in $r$ compared with the best-known bound[65, Thm. 3.23]. The direct rank $r+2$ truncation of an EDM (i.e., the SVD-MDS) has been shown to be minimax sub-optimal in [53, Sec. IV]. Since the OS-MDS shares almost same empirical phase transition edge as the SVD-MDS[1], it suggests that the idea of using $\|\frac{1}{p}\mathcal{P}_{\Omega}\mathbf{D}^{\star}-\mathbf{D}^{\star}\|$ to construct spectral estimator might be overly conservative in practice.
Remark II.2.
Similar to classic results mentioned in Remark II.1, step size rule in Lemma II.2 depends on both sample complexity and the radius of RIC. We do not opt to optimize $c_{c}$; instead, $c_{I}$ is explicitly left as a tunable constant. Enlarging $c_{I}$ makes the convergence of APGD more fragile, consistent with numerical observations in Section III-B. Skeptical reader may have pointed out that step size used in Lemma II.2 appears overly conservative. In pre-conditioning free LRMC and HMC contexts, more aggressive step sizes of the order $\mathcal{O}(1/\kappa^{2})$[52], $\mathcal{O}(1/\kappa)$[50][66] have been proven feasible. The $\mathcal{O}(1/\kappa)$ rate upon $\mathcal{O}(1/\kappa^{2})$ is obtained by reproducing higher order term of $\|\Delta\|_{F}$ when establishing regularity condition[66, Lemma 16, 17]. However, the $\mu$, $r$ dependency here can not be trivially moved. The largest component in restricted smoothness part lies in (37), where we utilize Lemma C.5 to bound $|R_{4}|$, thus additional $cn\|\Delta\|_{2,\infty}^{2}$ term appears. We infer that Lemma C.5 is not sharp up to $\sqrt{n}$, but we cannot amend it by assuming standard incoherence alone. If this estimate can be shown to match the random graph lemma[36, Lemma 7.1], it would immediately yield the local convergence mechanism, i.e., the contraction speed measured by $\|\Delta\|_{F}$, when optimizing the s-stress function.
II-C Why APGD and Why not $\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}$
Due to the projection property of $\mathcal{P}_{\Omega}$ under Bernoulli sampling model, it is clear that $\mathcal{P}_{\Omega}^{*}\mathcal{P}_{\Omega}=\mathcal{P}_{\Omega}$. Obviously, the orthonormality of the sensing basis $\mathbf{e}_{i}\mathbf{e}_{j}^{T}$ ensures this equivalence. However, for the EDMC problem, the primal basis correlation matrix $\mathbf{H}_{\boldsymbol{\omega}}$ and its inverse $\mathbf{H}_{\boldsymbol{\omega}}^{-1}$ (the dual basis correlation matrix[48]) are not diagonal. Therefore, theoretical analysis of optimizing $\frac{1}{p^{2}}\|\mathcal{R}_{\Omega}(\mathbf{PP}^{T}-\mathbf{G}^{\star})\|_{F}^{2}$ and the local behavior of its corresponding gradient, as suggested by[64], should be considerably more obscure. It has been pointed out in[1][42, Lemma B.1] that $\frac{1}{p^{2}}\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}$ is a biased estimator for $\mathcal{I}$, since
$\frac{1}{p^{2}}\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}(\cdot)=\sum_{\boldsymbol{\alpha}\neq\boldsymbol{\beta}}\frac{1}{p}\delta_{\boldsymbol{\alpha}}\frac{1}{p}\delta_{\boldsymbol{\beta}}\langle\cdot,\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle\langle\boldsymbol{\nu}_{\boldsymbol{\alpha}},\boldsymbol{\nu}_{\boldsymbol{\beta}}\rangle\boldsymbol{\omega}_{\boldsymbol{\beta}}$
$+\sum_{\boldsymbol{\alpha}=\boldsymbol{\beta}}\frac{1}{p^{2}}\delta_{\boldsymbol{\alpha}}\|\boldsymbol{\nu}_{\boldsymbol{\alpha}}\|_{F}^{2}\langle\cdot,\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle\boldsymbol{\omega}_{\boldsymbol{\alpha}}=\frac{1}{p^{2}}\tilde{\mathcal{Q}}_{\Omega}^{*}\mathbf{H}^{-1}_{\boldsymbol{\omega}}\tilde{\mathcal{Q}}_{\Omega}(\cdot),$ (18)
where $\mathbf{H}^{-1}_{\boldsymbol{\omega}}]_{\boldsymbol{\alpha}\boldsymbol{\beta}}=\langle\boldsymbol{\nu}_{\boldsymbol{\alpha}},\boldsymbol{\nu}_{\boldsymbol{\beta}}\rangle$ denotes the $(\boldsymbol{\alpha},\boldsymbol{\beta})$ element in the correlation matrix of $\{\boldsymbol{\nu}_{\boldsymbol{\alpha}}\}_{\boldsymbol{\alpha}\in\mathbb{I}}$, and $\mathbf{H}^{-1}_{\boldsymbol{\omega}}]_{\boldsymbol{\alpha}\boldsymbol{\alpha}}=\frac{n^{2}-2n+2}{2n^{2}}$ for any $\boldsymbol{\alpha}$[23, Lemma A.6]. [42] then introduce the idea of diagonal de-biasing into (18), and construct the following sensing operator
$\frac{1}{p^{2}}$ $\mathcal{M}_{\Omega}(\cdot)=\underbrace{\frac{1}{p^{2}}\tilde{\mathcal{Q}}_{\Omega}^{*}\mathrm{Od}(\mathbf{H}^{-1}_{\boldsymbol{\omega}})\tilde{\mathcal{Q}}_{\Omega}(\cdot)}_{\mathcal{M}^{1}_{\Omega}(\cdot)}+\underbrace{\frac{1}{p}\|\boldsymbol{\nu}_{\boldsymbol{\alpha}}\|_{F}^{2}\tilde{\mathcal{Q}}_{\Omega}^{*}\tilde{\mathcal{Q}}_{\Omega}(\cdot)}_{M_{2}}.$
Two-sided concentration property of $M_{2}$ was established in[22, Lemma B.1] and further refined by[42, Lemma B.6]. While deriving analogies to Lemma C.3 and C.5 for $\mathcal{M}^{1}_{\Omega}(\cdot)$, either in terms of tangent space Restricted Isometry Property (RIP) or non-tangent space uniform bounds, is challenging and unsatisfactory due to at least two reasons.
First, the desired estimates should hold uniformly over the set of low-rank matrices with nice incoherence[67, Thm. 22], in other words
$Z_{\mathcal{H}}(\boldsymbol{\xi}):$ $=\sup_{\mathbf{X}\in\mathbb{T},\,\|\mathbf{X}\|_{F}=1}\bigg|\langle\mathbf{X},(\frac{1}{p^{2}}\mathcal{M}^{1}_{\Omega}-\mathbb{E}\frac{1}{p^{2}}\mathcal{M}^{1}_{\Omega})(\mathbf{X})\rangle\bigg|$
$=\sup_{\mathbf{H}\in\mathcal{H}}|\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}^{T}-\mathbb{E}(\boldsymbol{\delta}^{T}\mathbf{H}\boldsymbol{\delta}^{T})|,$
will be considered, where
$\boldsymbol{\delta}:=[\delta_{(1,2)}/p,\dots,\delta_{(n-1,n)}/p]\in\mathbb{R}^{L},$
$\mathbf{H}:=\mathbf{W}_{\mathbf{X}}\circ\mathrm{Od}(\mathbf{H}^{-1}_{\boldsymbol{\omega}}),\,[\mathbf{W}_{\mathbf{X}}]_{\boldsymbol{\alpha}\boldsymbol{\beta}}:=\langle\mathbf{X},\boldsymbol{\omega}_{\boldsymbol{\alpha}}\rangle\langle\mathbf{X},\boldsymbol{\omega}_{\boldsymbol{\beta}}\rangle,$
$\mathcal{H}:=\{\mathbf{W}_{\mathbf{X}}\circ\mathrm{Od}(\mathbf{H}^{-1}_{\boldsymbol{\omega}}):\mathbf{X}\in\mathbb{T},\|\mathbf{X}\|_{F}\leq 1\}.$
Naively combining the $\epsilon$-net covering argument on set $\mathcal{H}$ with Hanson-Wright inequality[68, Ch. 6] causes the tail to blow up. In literature, similar structures are often resolved by the “suprema of chaos” inequality[69] for sub-Gaussian chaos generated by certain set of PSD matrices, which requires bounding Talagrand’s $\gamma_{2}$-functional[68, Ch. 8.5], e.g., [67, Sec. VI-B]. While we conjecture that an analogous tangent space eigenvalue lower bound as in[20, Lemma 10] for $\frac{1}{p^{2}}\mathcal{R}_{\Omega}^{*}\mathcal{R}_{\Omega}$ can be obtained using this technique, the de-biasing step causes $\mathbf{H}$ matrix no longer PSD, precluding a direct application of Krahmer’s result[69, Thm. 1.4]. Second, it is known that “suprema of chaos” based non-tangent space upper bound over residual with bounded $l_{2,\infty}$ or $l_{\infty}$ norm[70, Lemma 7, 8] does not guarantee finite-sample exact recovery for reasons other than sample splitting. Unlike Lemma II.2, both the perturbation error and sample complexity are related to the Frobenius norm of the total residual during each step of contraction induction argument, see [70, Sec. I-E] for further discussion.
Finding alternatives to bypass the aforementioned roadblocks is beyond the scope of this work. But it is interesting to point out that $\frac{1}{p^{2}}\mathcal{M}_{\Omega}$ appears to exhibit better numerical stability than $\frac{1}{p}\mathcal{R}_{\Omega}$, as empirically reported in [42, Sec. 8]. We frame this as a choice between analytical tractability and algorithmic fidelity, since our choice provides only marginal practical significance, whereas explaining the algorithm proposed in [42] requires substantially more effort.
III Numerical Experiments
The algorithm tested here is slightly different from the theoretical routine listed in Theorem II.1 (cf. Algorithm 2), in which BB stepsize without line search is employed, as inherited from[61, Ch. 2]. The performance of incoherent projection-free OS-MDS, OS-MDS initialized GD refinement on s-stress (OS-MDS-GD)[1], and OS-MDS initialized IFHT-EDMC (OS-MDS-IFHT)[23, Alg. 2] are also reported in Fig. 1 and Fig. 2(e), respectively.
III-A Implementation Details
where Lemma C.6 guarantees the approximation behavior of $\hat{\sigma}_{1}^{\star}$, and $c_{1}$, $c_{2}$ are the parameters to tune. Since the APGD is more analogous to an analytical process rather than an algorithm used for benchmarking, we assume throughout all experiments that $\sigma_{1}^{\star}$ and $\mu$ are perfectly known. Unless otherwise specified, the ground truth point set is generated by a standard Gaussian distribution with embedding dimension $r=2$, and results presented here are obtained under $\eta_{\max}=10$ [^11] The lower safeguard on BB stepsize is removed in most experiments, i.e., $\eta_{k}$ is allowed to be negative. Compared with pure GD, we found such sporadically occurring gradient ascent can numerically improve overall global convergence in all tested scenarios (especially the protein tests in Section III-E). The upper safeguard is retained to ensure algorithmic stability when $c_{IP}$ is large, c.f., (20).. Algorithm 2 is terminated once $\|\mathbf{p}g^{k}\|_{F}\leq 1\times 10^{-6}$ or the iteration number reaches $N=1000$. In contrast, both the upper and lower safeguards, $\eta_{\max}=\frac{2}{p\sigma_{1}^{\star}\mu r\kappa}$ and $\eta_{\min}=\frac{1}{2p\sigma_{1}^{\star}\mu r\kappa}$, are adopted in Section III-D for better illustration of linear convergence. Please note that the calculation of $\mathbf{JAJ}$ for $\mathbf{A}\in\mathbb{R}^{n\times n}$ can be conducted in $\mathcal{O}(n^{2})$ flops, as it amounts to sequentially removing the mean value of $\mathbf{A}$’s columns and rows.
Algorithm 2 APGD With $B$ Stepsize
@@ III-A Implementation Details
where $\mathrm{g}_{2}^{k}$ can be viewed as the norm of the second term in the gradient of (7a), it cannot be computed unless direct access to $\mathbf{G}^{\star}$ is available. We conjecture that the limited numerical performance of APGD in the sample-limited regime can be attributed to at least three factors: (i) the weakening of implicit regularization; (ii) the large residual between the pseudo-gradient and its population-level counterpart; (iii) the inconsistency between pseudo-gradient and the true gradient of (7a). These factors will be illustrated sequentially in the following sections.
Figure 1: The phase transition of regularization free OS-MDS and OS-MDS-GD when varying $p$ and nn[1]. We claim a success if the spectral error falls below 11 in (a), or the EDM recover rate is smaller than $10^{-3}$ in (b). For each $(n,p)$, the result is obtained by 100 independent trials.
Figure 2: (a) The phase transition of APGD when varying Bernoulli sample rate $p$ and the number of points $n$. The result is obtained by 100 independent trials. To verify Theorem II.1, we claim a success if the EDM recover rate falls below $10^{-3}$. In (b), (c), and (d), we vary the constant $c_{IP}$ in (20) when performing incoherent projection, while all other parameters remain unchanged. (e) depicts the phase transition of IFHT-EDMC algorithm [23, Alg. 2] under the same problem setup as in (b). Subplots (f), (g), and (h) show selected phase transition curves when $n=500,1000,1500$ with $c_{IP}$ varies.
@@
Figure 4: Trajectories of all three quantities in (19) while fixing $n=1500$. The result is obtained from 100 independent trials. (a), (d) shows the tendency of $r^{k}$ in selected succeed instances (96% when $p=0.1$, 47% when $p=0.05$). (b), (c) plots the corresponding $g_{1}^{k}$, $g_{2}^{k}$ when $p=0.1$. (e), (f) contains the records of $g_{1}^{k}$ and $g_{2}^{k}$ of selected failed instances (53% when $p=0.05$). These trajectories can be divided into two groups (17% in region 1, 36% in region 2).
III-B Phase Transition Under Gaussian Point Set
in (16) with some $c_{IP}>1$, the performance of APGD on Gaussian random point sets will be degenerated, as suggested by Fig. 2(b), (c), and (d). This observation indicates that the power of implicit regularization to maintain incoherence is weakened in the APGD case. More delicate evaluations are carried out in Fig. 2(e), (f), and (g). This reveals that APGD can still converge when the trimming step is removed, yet larger $c_{IP}$ gives rise to more oscillations and non-monotonicity in the transition curve, ultimately leading to broader transition interval measured by $p$. Such conclusion is also supported by Fig. 2(e), where the OS-MDS-IFHT[23], i.e., regularization-free embedding geometry counterpart of APGD, exhibits a resemblance in its transition edge to that in Fig. 2(d)  [^12] Since IFHT is a tricky variant of Singular Value Projection algorithm[71], which eliminates the necessity of proving the iteration can automatically stay incoherent[72, Sec. III-B]..
By comparing Fig. 1(a) and Fig. 2(a), we find that the phase transition edge of APGD matches that of OS-MDS. This phenomenon confirms the prediction in Lemma II.2, and yet is quite unusual, as both non-convex Gaussian ensemble phase retrieval and structure-less LRMC do not necessarily rely on the spectral initialization to succeed[73][74]. Since the proof roadmap in[74] indicates that the success of rescaled eigenspace alignment initialization relies on the implicit incoherent regularization ability of the original fourth-order polynomial-type cost function for LRMC, we infer that when this property is compromised, the algorithm may start to behave selectively on the quality of starting point. Alternatively, more dedicated initialization for APGD can somewhat improve the transition edge. However, the OS-MDS-GD strategy exhibits empirical convergence when started inside a slack region around the ground truth as depicted in Fig. 1(b), even without any incoherence projection.
TABLE II: Parameters of Test Proteins
@@ III-C Phase Transition Under Random Perturbation
where $\mathbf{E}_{n}$ is a Gaussian random matrix with $\mathbf{E}_{n}]_{ij}\sim\mathcal{N}(0,1)$ and $\mathbf{P}^{\star}$ is generated from a uniform distribution over the hypercube $0.5,0.5]^{r}$. The noise level $\sigma_{n}$ is varied from $5\times 10^{-5}$ to $10$. The iteration will be terminated if $\|\mathbf{p}_{g}^{k}\|_{F}\leq 10^{-8}$, while all other parameters follow the same setup as in Fig. 2(b). Since the perturbation can be relatively small, we claim a success if $\mathrm{RE}\leq 10^{-5}$, and the resulting phase transitions are plotted in Fig. 3 for $n=500,1000,1500$, $r=2$, where $\|\Delta_{0}\|_{F}^{2}$ refers to $\mathrm{dist}(\mathbf{P}_{0},\mathbf{P}^{\star})^{2}$. Around $\sigma_{n}\leq 0.1$, i.e., inside the region where $\|\Delta_{0}\|_{F}^{2}\leq\sigma_{r}^{\star}$ according to Fig. 3(d), these plots depict that the convergence of APGD becomes sensitive primarily to the sample complexity. As $p\to 1$, the $\hat{\nabla}f(\mathbf{P})$ reduces to the gradient direction of the vanilla matrix factorization problem, which is known to exhibit local restricted strong convexity[75, Thm. 5]. Therefore, Fig. 3 also demonstrates the lower bound on $p$ necessary for this approximation to hold.
However, this dependence on $p$ is sub-optimal compared to the performance achieved by either OS-MDS-GD or trace minimization methods[20, Sec. IV][31, Sec. 5], indicating that the stabilization of pseudo gradient, e.g., ensuring the upper bound on Eq. (26) in Appendix B-B, needs more random samples to guarantee, even though Lemma II.2 predicts that the establishment of regularity condition should require the same order of sample complexity as in classical LRMC problems. Moreover, it is noteworthy that OS-MDS cannot achieve $\|\Delta_{0}\|_{F}^{2}\leq\sigma_{r}^{\star}$ with high probability at the sample configurations $\{(n,p):(500,0.08),(1000,0.05),(1500,0.04)\}$, which is why we allow Algorithm 2 to take negative stepsize, thereby facilitating global optimization.